# Data-Effective Learning: A Comprehensive Medical Benchmark

### v1

@@ 3.3 The benchmark evaluation metric: NormDEL
(a) Box-plot of Kvasir-Instrument
(b) Box-plot of Kvasir-SEG
(a) Box-plot of Kvasir-Instrument
@@
(e) Box-plot of CVC-ClinicDB
(d) Box-plot of ETIS
(e) Box-plot of CVC-ClinicDB
(f) Box-plot of CVC-ColonDB
Figure 4: Box plot of the baseline method (MedDEL) for testing mIoU under retaining different data ratios. Under the same computational resources, we plot the impact of varying proportions of pre-training data on downstream task performance, using the data proportions from Table 2 and models pre-trained on ImageNet for reference. The results demonstrate that the performance of downstream tasks initially improves with an increase in pre-training data but eventually decreases. This is because the initially added data is effective, however, as more data is introduced beyond a certain point, it includes redundancy or erroneous information, which adversely affects the model’s performance.
4 Experiment
@@
4.1 Settings
To achieve the pretraining of the Foundation Model, enabling it to provide a better initial value for the majority of endoscopy tasks, we use the unsupervised Masked Auto Encoder (MAE-ViT-Base) He et al. (2021a).
During the training of all models, we selectively filter the pretraining dataset capacity based on different thresholds of $\epsilon$ and $\eta$. Specifically, we set $\epsilon$ to 0.9 and $\eta$ from 0.7 to 0.9 with a step size of 0.05. Consequently, we obtain the dataset quantities as presented in Table 2. The batch size is set to 16, and the training is conducted using AdamW Loshchilov and Hutter (2017) and a cosine learning rate Gotmare et al. (2018). The model maintains a peak learning rate of $1.5\times 10^{-4}$ throughout the training process. In the validation of downstream tasks, we implement a simple Multi-Layer Perceptron (MLP) Tolstikhin et al. (2021) to achieve segmentation heads.
4.2 Effective Data Utilization with MedDEL
Dataset Pretraining Time mIoU 20% pretraining data 33% pretraining data
@@ 4.2 Effective Data Utilization with MedDEL
Table 4: Illustrating the impact of different pretraining time (epochs) on performance metrics (mIoU) using 20% and 33% of pretraining data. The results indicate that the optimal performance is often not achieved with the standard pretraining duration (1000 epochs for 20% and 600 epochs for 33%) but rather occurs with early pretraining time, demonstrating the effectiveness of MedDEL in reducing computational resource consumption.
Dataset mIoU 5% pretraining data 20% pretraining data 33% pretraining data 50% pretraining data
MedDEL Random MedDEL Random MedDEL Random MedDEL Random
CVC-300 63.81 60.17 61.40 60.53 62.84 60.27 62.60 62.19
CVC-ColonDB 69.46 68.62 69.85 69.01 71.54 69.50 69.71 69.48
Kvasir-SEG 75.67 75.33 76.23 76.04 75.96 75.75 76.60 76.13
PolypGen2021 60.92 60.70 61.51 61.47 62.24 60.46 62.19 60.91
Kvasir-Instrument 79.50 79.24 80.11 79.67 80.29 79.24 80.35 79.23
Table 5: MedDEL vs. random selection method. For both MedDEL and random method, the influence of various proportions of pretraining data on the average mIoU over the last 50 epochs was investigated across 5 different datasets. MedDEL outperforms random method across any proportion of the pretraining dataset. Moreover, as the size of the pretraining dataset decreases, the average mIoU over the last 50 epochs does not exhibit a significant decrease. This underscores the effectiveness of utilizing MedDEL for dataset structural optimization.
4.3 Storage Saving Analysis
@@
4.4 Computing Power Saving Analysis
@@
4.5 Ablation Study
To validate the effectiveness of our model, we conduct ablation experiments where we compare the MedDEL method and the random method across different downstream task datasets. The metric utilized in Table 5 is the average mIoU over the last 50 epochs, chosen to mitigate errors compared to relying solely on the mIoU from the last epoch. The results are promising, consistently showcasing our approach outperforming random selection.
Specifically, on the CVC-300, CVC-ColonDB, Kvasir-SEG, PolypGen as well as Kvasir-Instrument datasets, MedDEL’s performance significantly surpasses that of random selection method with all different proportions of pretraining data. This is particularly evident in the more favorable mIoU (mean intersection over union) curves observed throughout the entire training process, which clearly demonstrates the stability and excellence of our approach.
5 Conclusion
We propose the first medical data-effective Benchmark, which involves a million-level dataset from 31 medical centers (DataDEL), a data-effective approach (MedDEL), and a comprehensive metric for evaluating pretraining data volume (NormDEL). The establishment of an open benchmark for data-effective learning is crucial for the medical artificial intelligence research community, which is an encouraging foundation for subsequent research endeavors.

### v3

@@ 3.3. The benchmark evaluation metric: NormDEL
(a) Box-plot of Kvasir-INS
@@
(e) Box-plot of PolyGen2021
(f) Box-plot of CVC-300
(g) Box-plot of CVC-ClinicDB
(h) Box-plot of CVC-ColonDB
Figure 4. Box plot of the baseline method (MedDEL) for testing mIoU under retaining different data ratios. Under the same computational resources, we plot the impact of varying proportions of pre-training data on downstream task performance, using the data proportions from Table 2 and models pre-trained on ImageNet for reference. The results demonstrate that the performance of downstream tasks initially improves with an increase in pre-training data but eventually decreases. This is because the initially added data is effective, however, as more data is introduced beyond a certain point, it includes redundancy or erroneous information, which adversely affects the model’s performance.
4. Experiment
@@
4.1. Settings
To achieve the pretraining of the Foundation Model, enabling it to provide a better initial value for the majority of endoscopy tasks, we use the unsupervised Masked Auto Encoder (MAE-ViT-Base) (He et al., 2021a).
During the training of all models, we selectively filter the pretraining dataset capacity based on different thresholds of $\epsilon$ and $\eta$. Specifically, we set $\epsilon$ to 0.9 and $\eta$ from 0.7 to 0.9 with a step size of 0.05. Consequently, we obtain the dataset quantities as presented in Table 2. The batch size is set to 16, and the training is conducted using AdamW (Loshchilov and Hutter, 2017) and a cosine learning rate (Gotmare et al., 2018). The model maintains a peak learning rate of $1.5\times 10^{-4}$ throughout the training process. In the validation of downstream tasks, we implement a simple Multi-Layer Perceptron (MLP) (Tolstikhin et al., 2021) to achieve segmentation heads.
Table 4. Illustrating the impact of different pre-training time (epochs) on performance metrics (mIoU) using 20% and 33% of pretraining data. The results indicate that the optimal performance is often not achieved with the standard pretraining duration (1,000 epochs for 20% and 600 epochs for 33%) but rather occurs with early pretraining time, demonstrating the effectiveness of MedDEL in reducing computational resource consumption.
Dataset Pretraining Time 20% pretraining data 33% pretraining data
@@ 4.1. Settings
4.2. Effective Data Utilization with MedDEL
Figure 5. Demonstration of images deleted by MedDEL. This figure shows MedDEL deleting semantically similar images, which appear to have no significant differences between them from a perceptual perspective,
Figure 6. Model generalization experiments across different datasets with different data volumes. The experiments included four distinct datasets: Kvasir-SEG, ImageCLEFmed, CVC-ClinicDB, and CVC-ColonDB. The results indicate the performance of the model at different volumes of pre-training data (5%, 20%, 50%, and 100%), and compared the outcomes between randomly selected data and data selected using the MedDEL method.
4.3. Storage Saving Analysis
@@
4.4. Computing Power Saving Analysis
@@
4.5. Visualization of MedDEL(DEL)
The MedDEL algorithm, as a benchmark method in the field of endoscopic, is based on the Vision Transformer (ViT) model  (Dosovitskiy et al., 2020) for extracting image features, generating a 768-dimensional feature output. By introducing K-means clustering  (Macqueen, 1967), it successfully reduces computational complexity and further effectively addresses irrelevant data in the endoscopic dataset by setting similarity thresholds and filtering conditions.
Figure 5 visualizes the results of the MedDEL (DEL) algorithm, where upon observation, it is evident that MedDEL eliminates a significant portion of redundant data with highly similar features from the dataset. Considering their similarity, we suppose removing such redundant data is a reasonable strategy to optimize the dataset structure, enhancing dataset quality. This optimization improves the model’s understanding of endoscopic images, making it more targeted and interpretable, and providing robust support for medical image analysis  (Dosovitskiy et al., 2020).
4.6. Ablation Study
To validate the effectiveness of our model, we supplement the following generalization experiments in Fig 6. Specifically, we select 5%, 20%, and 50% of the data and compare the test performance of random selection and MedDEL. We find that whether using the random selection method or MedDEL, as the amount of data increases, the model’s generalization ability improves. Furthermore, at the same data volume, MedDEL shows superior performance to the random method. Additionally, MedDEL shows comparable generalization performance when selecting 50% data compared to using full data.
5. Conclusion
In the era of foundation models, we are the first to propose whether larger pre-training data necessarily leads to improved model performance. In order to investigate this issue, we propose the first medical data-effective Benchmark, which involves a million-level dataset from 31 medical centers (DataDEL), a data-effective approach (MedDEL), and a comprehensive metric for evaluating pretraining data volume (NormDEL). The establishment of an open benchmark for data-effective learning is crucial for the medical artificial intelligence research community, which is an encouraging foundation for subsequent research endeavors.
Acknowledgement
This work was supported in part by NSFC (No. U2001209, 62372117). The computations in this research were performed using the CFFF platform of Fudan University.